The Caverns of Sonora have two different types of tours: the Crystal Palace tour and the Horseshoe lake tour. The total length of both tours 3.25 mi. The Crystal Palace tour is a half mile less than twice the distance of Horseshoe Lake tour. Determine the length of each tour.

Answer by macston(5194)

.
C=length Crystal Palace tour; H=length Horseshoe Lake tour
.
C=2H-0.5
.
C+H=3.25
(2H-0.5)+H=3.25
3H=3.75
H=1.25
ANSWER 1: Horseshoe Lake tour is 1.25 miles
.
C=2H-0.5=2(1.25)-0.5=2.5-0.5=2
ANSWER: Crystal Palace tour is 2.0 miles.
.